OpenStudy (austinl):

Parametric and Polar Curves My question requires that I convert polar equations to Cartesian Coordinates. And then Describe the resulting curve. I am given \[r=3csc(\theta)\]

terenzreignz (terenzreignz):

Another amusement?

terenzreignz (terenzreignz):

:D

OpenStudy (austinl):

Unfortunately no.... D:

terenzreignz (terenzreignz):

Well, this is a horizontal line :D

OpenStudy (austinl):

I know that I need to multiply each side by r. \[r^2=3r\csc(\theta)\]

OpenStudy (austinl):

\[r^2=x^2+y^2\]

terenzreignz (terenzreignz):

Forget that, \[\Large r = \frac3{\sin(\theta)}\]\[\Large r\sin(\theta) = 3\]

terenzreignz (terenzreignz):

hehe \[\Large \sin(\theta) = \frac{y}r\]

terenzreignz (terenzreignz):

Everything okay? :)

OpenStudy (austinl):

I'm trying do decipher lol.

terenzreignz (terenzreignz):

Sorry... \[\Large \sin(\theta) = \frac{y}r\]\[\Large y = r\sin(\theta)\] That said... \[\Large r = 3\csc(\theta)\]\[\Large r\sin(\theta) = 3\]\[\huge y = 3\]

OpenStudy (austinl):

As well, I am trying to figure out why the example from the book is ZERO HELP!!!!!!

terenzreignz (terenzreignz):

Zero?

OpenStudy (austinl):

The book example is \[r=6\sin(\theta)\] \[r^2=6r\sin(\theta)\] \[x^2+y^2=6r\sin(\theta)\] AND EVIDENTLY \[y=r\sin(\theta)\] which gives \[x^2+y^2-6y=0\] \[x^2+(y-3)^2=9\] Circle radius 3, centered at (0,3)

OpenStudy (austinl):

I mean, where do they get this arbitrary equation \[y=r\sin(\theta)\]

terenzreignz (terenzreignz):

That's not arbitrary :) It's Polar coordinates :D Say that is your point (x,y)

terenzreignz (terenzreignz):



terenzreignz (terenzreignz):

So clearly, \[\large \sin \theta = \frac{y}r\] r represents the distance of said point from the origin .

OpenStudy (austinl):

AAAAHAAAAA! y u no explain this to me book!!!??!?!?!?!?//1/!?

terenzreignz (terenzreignz):

Book is arrogant :3

OpenStudy (austinl):

It really is a very poorly written book.

terenzreignz (terenzreignz):

LOL Issue a protest :D

OpenStudy (austinl):

Tried that....

terenzreignz (terenzreignz):

OH well, that's what OS is for :3

terenzreignz (terenzreignz):

err, dealing with the arrogance of said books :D

OpenStudy (austinl):

Thank you so much. It gives one example and expects us to apply it to tons of other problems.

OpenStudy (austinl):

Would you care to explain another one? I think I have an idea but the concept is still rough.

terenzreignz (terenzreignz):

Sure :)

OpenStudy (austinl):

I shall post another question so that you may be justly rewarded.